OpenStudy (anonymous):

The perimeter of a rectangle is 42, and its diagonal is 15. Find its length and width. length _____(larger side) width ______(smaller side)

OpenStudy (anonymous):

Rectangle perimiter: 2(a+b)=42 ; a+b=21 ; b=21-a Pythagoras: a^2+b^2=15^2 a^2+(21-a)^2-225=0 a^2-21a+108=0 Two solutions: a1 = 9 ; b=21-9 = 12 a2=12 ; b=21-12 = 9 So length=12, width=9